# The Thurston norm of graph manifolds

### v1

@@ 3 The Thurston norm of a graph manifold
For every $i=1,...,n$, let $\psi_{i}:H_{2}(M;\operatorname{\mathbb{Z}})\to\operatorname{\mathbb{Z}}$ be the map that associates to the class of a properly embedded oriented surface $S\subset M$ the algebraic intersection $l_{i}=\psi_{i}([S])$ between $S$ and a regular fiber of $M_{i}$. Extend $\psi_{i}$ linearly to a map $H_{2}(M;\operatorname{\mathbb{Q}})\to\operatorname{\mathbb{Q}}$ and then continuously to a linear map $H_{2}(M)\to\operatorname{\mathbb{R}}$. Finally, define the linear map $\psi=(\psi_{1},...,\psi_{n}):H_{2}(M)\to\operatorname{\mathbb{R}}^{n}$.
@@
where $x_{M_{i}}$ is the Thurston norm on $H_{2}(M_{i},\partial M_{i})$. Each $M_{i}$ is a Seifert fibered manifold and the proof of Proposition 2.14 shows that
$x_{M_{i}}([S\cap M_{i}])=-\chi_{i}|\psi_{i}([S])|.$
@@
where $V$ is the null-space of the Thurston norm on $H_{2}(M)$. In particular, the nonvanishing homology $H_{2}^{nv}(M)$ is isomorphic to $\ker A\subset\operatorname{\mathbb{R}}^{n}$ through an isomorphism sending $H_{2}^{nv}(M;\operatorname{\mathbb{Q}})$ to $\ker A\cap\operatorname{\mathbb{Q}}^{n}$.
@@
is exact. Since $\chi_{i}<0$ for every $i$, $M$ cannot be a torus-bundle. Indeed, a non-Seifert fibered torus-bundle has JSJ decomposition equal to a product (torus)×[0,1]$\times$[0,1]=× [ 0 , 1 ] =(annulus)×$\times S^{1}$, given by cutting along a torus fiber. Hence, Lemma 2.17 implies that $\psi(H_{2}(M;\operatorname{\mathbb{Q}}))=\ker A\cap\operatorname{\mathbb{Q}}^{n}$. Finally, the equality (1) implies $V\cap H_{2}(M,\operatorname{\mathbb{Q}})=\ker\psi$. ∎
@@
The statement is a direct consequence of formula (1) after identifying $\operatorname{Im}\psi=\ker A$. ∎
Proposition 3.3 expresses $x^{nv}$ as a sum of absolute values of linear functionals. Construction 5.9 describes the polyhedron of such a norm.
@@
Thanks to the previous proposition, if $S\subset M$ is an oriented and properly embedded surface of norm zero, then $S\cap M_{i}$ is isotopic to a vertical surface in $M_{i}$ for every $i$. Since the regular fibers of two adjacent pieces of $M$ are never parallel, $S$ is a disjoint union of closed vertical surfaces in some $M_{i}$s. For each $i$, let $\overline{V_{i}}\subset H_{2}(M_{i},\partial M_{i})$ be the subspace generated by the classes of vertical closed surfaces in $M_{i}$. Choose a subspace $V_{i}\subset H_{2}(M)$ such that the map $H_{2}(M_{i})\to H_{2}(M_{i},\partial M_{i})$ restricts to an isomorphism between $V_{i}$ and $\overline{V_{i}}$. Identify $V_{i}$ with its image in $H_{2}(M)$. Let $\mathcal{T}\subset H_{2}(M)$ be the subspace generated by the classes of JSJ tori of $M$. The null-space of the Thurston norm splits as the direct sum $\mathcal{T}\oplus V_{1}\oplus...\oplus V_{n}$. If we show that the dimension of $\mathcal{T}$ is $b_{1}(\Gamma)$, then the statement follows from Propositions 2.15 and 3.2.
@@
We first show that a class $\alpha\in H_{2}(M;\operatorname{\mathbb{Z}})$ is fibered if and only if $\psi_{i}(\alpha)\neq 0$ for each $i=1,...,n$. Suppose that $\psi_{i}(\alpha)\neq 0$ for each $i$. Without loss of generality, assume that $M$ is connected and $\alpha$ is primitive, so that $\alpha$ can be represented by a connected properly embedded oriented surface $S\subset M$. We can also assume that $S\cap M_{i}$ is a (possibly disconnected) horizontal surface in $M_{i}$. Cut $M$ open along $S$, then call $N$ the new manifold. Notice that the manifold $N_{i}$ obtained by cutting $M_{i}$ open along $S\cap M_{i}$ is an oriented interval-bundle over $S\cap M_{i}$, hence necessarily $(S\cap M_{i})\times[0,1]$. Properly speaking, $N_{i}$ is a manifold with corners, $\partial N_{i}$ being the union of the “vertical boundary” $(S\cap\partial M_{i})\times[0,1]$ and the “horizontal boundary” $(S\cap M_{i})\times\{0,1\}$. Since $\alpha\neq 0$, $S$ is non-separating and $N$ is the connected manifold obtained by gluing the interval bundles $(S\cap M_{i})\times[0,1]$ along their “vertical” boundary components. Each vertical boundary component is an interval-foliated annulus. Since the foliations of the annulus by intervals are all isotopic with each other, we can slightly isotope the interval-bundles of each $N_{i}$ to get an interval-bundle of $N$ over the surface $S$. Again, this interval-bundle must be $S\times[0,1]$, thus $M$ fibers over the circle with fiber $S$.
For the vice-versa, suppose that $M$ fibers over the circle with fiber $S$. As in the proof of Proposition 3.2, $S$ cannot be a torus. In particular, since the fibration is a taut foliation, by [Th86] each JSJ torus of $M$ can be slightly isotoped to be transverse to the fibers of the fibration, apart for a finite number of saddle singularities. In fact, by the Poincare-Hopf Theorem, each JSJ torus can be taken to be entirely transverse to the fibration. Hence, the fibration restricts to a fibration onto every piece $M_{i}$. If any $S\cap M_{i}$ were vertical, then $M_{i}$ would be a torus-bundle or an annulus-bundle, which is excluded since $\chi_{i}<0$. It follows that $\psi_{i}([S])\neq 0$, for every $i=1,...,n$.
@@ 4 Characterization of the norms that can be realised
The formula (1) implies that the nonvanishing Thurston norm of a closed oriented graph manifold can be expressed as a sum of absolute values of linear functionals $H_{2}^{nv}(M)\to\operatorname{\mathbb{R}}$. Notice that the linear functionals are rational on $H_{2}^{nv}(M;\operatorname{\mathbb{Q}})$. In this section, we will prove that every norm with these properties can be realised as the nonvanishing Thurston norm on some closed oriented good graph manifold.
@@
Let $\beta_{1},...,\beta_{n}\in\operatorname{\mathbb{Q}}^{d}$ be such that the function
@@
is a norm on $\operatorname{\mathbb{R}}^{d}$, and let $g_{1},...,g_{n}$ be nonnegative integers. There exist a closed oriented good graph manifold $M$ with base surfaces of genera $g_{1},...,g_{n}$ and a rational basis on $H_{2}(M)$, such that the nonvanishing Thurston norm on $H_{2}^{nv}(M)$ in the induced basis coincides with |||||$\cdot$||| | ⋅ | |. Moreover, we can ask for either all or none of the top-dimensional Thurston cones of $H_{2}(M)$ to be fibered.
@@
Finally, for the fibering statement, recall that a class $\alpha\in H_{2}(M)$ is fibered if and only if $\psi_{i}(\alpha)\neq 0$ for each $i=1,...,n$ (see the proof of Proposition 3.5). In particular, if $M$ does not fiber over the circle, there is an $i$ so that $\psi_{i}(\alpha)=0$ for every $\alpha\in H_{2}(M)$. Hence, if we want $M$ to fiber over the circle, then we just need to take $\beta_{1},\beta_{2},...,\beta_{n}$ all nonzero. Oppositely, if we want $M$ not to fiber over the circle, then it is enough to ask for one of the $\beta_{1},...,\beta_{n}$ to be zero (possibly by using $n+1$ instead of $n$ vectors). ∎
@@
The same reasoning as the one in [Th86, Chapter 4] shows that every such norm is a sum of absolute values of linear functionals $\operatorname{\mathbb{Q}}^{2}\to\operatorname{\mathbb{Q}}$. The statement follows from Theorem 4.1 by choosing all $g_{i}=0$. ∎
Unfortunately, we will see in the next section that the analogous statement does not hold in higher dimension, and that the Thurston norms of graph manifolds constitute a strictly smaller family of the Thurston norms of $3$-manifolds.
@@ 5 Linear cellular decompositions of the sphere induced by Thurston norms
Now, we aim to explore geometrically the nonvanishing Thurston norms of graph manifolds. We will describe how to recover the shape of their unit polyhedra and observe a key property of them, that we will call “completeness”. Looking for completeness of the unit balls will allow us to show that the family of norms realisable as nonvanishing Thurston norms of graph manifolds is strictly contained in the family of nonvanishing Thurston norms of compact oriented $3$-manifolds.
@@
Let $P\subset\operatorname{\mathbb{R}}^{n}$ be a convex (solid) polyhedron containing the origin in its interior. Let $\pi:\operatorname{\mathbb{R}}^{n}-\{0\}\to S^{n-1}$ be the radial projection. The cellular decomposition of $\partial P$ in faces induces a cellular decomposition of $S^{n-1}$ via the projection $\pi$. If $F$ is a $k$-dimensional (open) face of $\partial P$, then $\pi(F)$ is a $k$-dimensional cell in $S^{n-1}$, contained in the intersection of $S^{n-1}$ with a $(k+1)$-dimensional subspace of $\operatorname{\mathbb{R}}^{n}$, for every $k\leq n-1$. We call such a cellular decomposition of $S^{n-1}$ a linear cellular decomposition and we indicate it with the symbol $\mathcal{F}(P)$. For $k=0,...,n-1$, $\mathcal{F}^{k}(P)$ will indicate the $k$-skeleton of $\mathcal{F}(P)$.
@@
When the inclusion above is an equality, we say that the linear cellular decomposition is complete, and the polyhedron $P$ is complete. When the $\beta_{i}$ are all in $\operatorname{\mathbb{Q}}^{n}$, we refer to the linear cellular decomposition as a rational cellular decomposition.
If |||||$\cdot$||| | ⋅ | | is a norm on $\operatorname{\mathbb{R}}^{n}$ whose unit ball in a polyhedron $P$, we call $\mathcal{F}(||\cdot||):=\mathcal{F}(P)$. Such a linear cellular decomposition is symmetric, meaning that $\mathcal{F}^{k}(||\cdot||)$ is preserved by the antipodal map $v\to-v$, for every $k=0,...,n-1$.
@@
Let $C$ be the cube $[-1,1]^{3}\subset\operatorname{\mathbb{R}}^{3}$, see Figure 2 left. The linear cellular decomposition of $S^{2}$ induced by $C$ is rational but not complete, indeed
@@
The same line of reasoning as above shows that $v$ belong to the open cone $C^{\prime}$ where $||\cdot||^{\prime}$ coincides with a linear function $L:\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}$. Since $C^{\prime}$ is open, there is an $\epsilon>0$ such that both $v+\epsilon\beta_{i}$ and $v-\epsilon\beta_{i}$ lie in $C^{\prime}$. We can then compute
@@
Let $L$ be a $3$-link chain and let $M:=S^{3}-\overset{\circ}{N}(L)$, see Figure 3. As shown in [Th86], the Thurston unit ball on $H_{2}(M,\partial M)$ is the polyhedron with vertices $(1,0,0),(0,1,0),(0,0,1),(-1,1,1)$ and their opposites. The linear cellular decomposition of $S^{2}$ induced by the Thurston norm is not complete, as there are vertices with odd degree. Hence, the family of norms that are realizable as nonvanishing Thurston norm of graph manifolds is strictly included in the family of norms realizable as nonvanishing Thurston norm of $3$-manifolds.
@@
In [Th86], Thurston showed that if a convex symmetric polygon in $\operatorname{\mathbb{R}}^{2}$ has vertices in $\operatorname{\mathbb{Z}}^{2}$, with entries accordingly congruent modulo $2$, then it can be realised as the dual Thurston ball of a compact orientable $3$-manifold. As pointed out by Sane in [Sa22], the idea at the base of Thurston’s proof is that every such polygon is the dual unit ball of an intersection norm on the first homology group of a torus. For every intersection norm on the first homology of a torus, Thurston finds a $3$-manifold whose Thurston norm on the second homology coincides with the given intersection norm. In particular, Sane generalizes this construction to the higher-dimensional case: he considers intersection norms on the first homology of the closed orientable surface of genus $g$ and their induced polyhedra in $\operatorname{\mathbb{R}}^{2g}$. As shown in [Sa20], not every polyhedron is realizable this time, nor with the same assumptions working for polygons.
@@
It is worth stressing that both Corollaries 5.7 and 5.8 are here established with respect to the choice of a rational basis for the second homology of the relative graph manifold. Observe that, in order to show the same results with respect to integral bases, a weaker version of the integer analogous to Theorem 4.1 would be enough. For instance, we would like that for every $\beta_{1},...,\beta_{n}$ integral generators of $\operatorname{\mathbb{Z}}^{d}$ there existed a closed oriented good graph manifold $M$, an integral basis of $H_{2}(M)$ and a natural $N$, such that the nonvanishing Thurston norm is $N\sum_{1}^{n}|\langle\beta_{i},\cdot\rangle|$ in the induced integral basis.
@@
$\operatorname{Span}_{\operatorname{\mathbb{R}}}\{\beta_{j}\,|\;j\neq i\}$ has dimension $n-1$, then choose $\overline{\beta}_{i}$ to be a non null vector in $\bigcap_{j\neq i}\beta_{j}^{\bot}$, normalized to have $||\overline{\beta}_{i}||=|\langle\overline{\beta}_{i},\overline{\beta}_{i}\rangle|=1$. The unit ball of |||||$\cdot$||| | ⋅ | | is the polyhedron with vertices ±$\pm\overline{\beta}_{1},...,\pm\overline{\beta}_{n}$ in $\operatorname{\mathbb{R}}^{n}$.
@@
Identity (4) was proved by Gabai in [Gab83]. In particular, the Thurston ball of $H_{2}(M)$ can be fully recovered by intersecting the Thurston ball of $H_{2}(\widetilde{M})$ with the subspace $\pi^{*}H_{2}(M)$ and rescaling it by a factor $2$.
@@
By Lemma 3.1 and Proposition 5.3, the nonvanishing Thurston unit ball on the nonvanishing second homology of a closed oriented graph manifold is a symmetric complete polyhedron, and it has rational vertices. Unfortunately, not every symmetric complete polyhedron with rational vertices can be obtained as the nonvanishing Thurston ball of some closed oriented graph manifold. As a counterexample, consider the polyhedron $P\subset\operatorname{\mathbb{R}}^{3}$ in Figure 5 with vertices

### v2

@@ 3 The Thurston norm of a graph manifold
For every $i=1,...,n$, let $\psi_{i}\in H^{2}(M)\cong\hom(H_{2}(M),\operatorname{\mathbb{R}})$ be the Poincare dual of a regular fiber $l_{i}\subset M_{i}$. Finally, define the linear map $\psi=(\psi_{1},...,\psi_{n}):H_{2}(M)\to\operatorname{\mathbb{R}}^{n}$.
@@
where $x_{M_{i}}$ is the Thurston norm on $H_{2}(M_{i},\partial M_{i})$. Each $M_{i}$ is a Seifert fibered manifold and equality (1) concludes the proof. ∎
@@
where $V$ is the null-space of the Thurston norm on $H_{2}(M)$. In particular, the nonvanishing homology $H_{2}^{nv}(M)$ is isomorphic to $\ker A\subset\operatorname{\mathbb{R}}^{n}$ through an isomorphism sending $H_{2}^{nv}(M;\operatorname{\mathbb{Q}})$ onto $\ker A\cap\operatorname{\mathbb{Q}}^{n}$. Hence, a rational basis of $H_{2}^{nv}(M)$ corresponds through $\psi$ with a rational basis of $\ker A$.
@@
is exact. Since $\chi_{i}<0$ for every $i$, $M$ cannot be a torus-bundle. Indeed, a non-Seifert fibered torus-bundle has JSJ decomposition equal to a product (torus)×[0,1]$\times$[0,1]=× [ 0 , 1 ] =(annulus)×$\times S^{1}$, given by cutting along a torus fiber. Hence, Lemma 2.17 implies that $\psi(H_{2}(M;\operatorname{\mathbb{Q}}))=\ker A\cap\operatorname{\mathbb{Q}}^{n}$. Finally, given that $\chi_{i}<0$ for each $i$, equality (4) implies $V\cap H_{2}(M,\operatorname{\mathbb{Q}})=\ker\psi$. ∎
@@
Let $\alpha_{1},...,\alpha_{d}$ be the basis of $H_{2}^{nv}(M)$ corresponding with $v_{1},...,v_{d}$ through $\psi$, i.e. such that $\psi(\alpha_{j})=v_{j}$ for every $j$.
For any $i=1,...,n$ and $a_{1},...,a_{d}\in\operatorname{\mathbb{R}}$,
$\psi_{i}(a_{1}\alpha_{1}+...+a_{d}\alpha_{d})=a_{1}\psi_{i}(\alpha_{1})+...+a_{d}\psi_{i}(\alpha_{d})=a_{1}(P)_{i1}+...+a_{d}(P)_{id}=\langle\beta_{i},(a_{1},...,a_{d})\rangle.$
Now, formula (4) implies the thesis. ∎
Proposition 3.3 expresses $x^{nv}$ as a sum of absolute values of linear functionals. Construction 5.9 describes the unit polyhedron of such a norm.
@@
Thanks to Lemma 3.1, if $S\subset M$ is an oriented and properly embedded surface of norm zero, then $S\cap M_{i}$ is isotopic to a vertical surface in $M_{i}$ for every $i$. Since the regular fibers of two adjacent pieces of $M$ are never parallel, $S$ is a disjoint union of closed vertical surfaces in some $M_{i}$s. For each $i$, let $\overline{V_{i}}\subset H_{2}(M_{i},\partial M_{i})$ be the subspace generated by the classes of vertical closed surfaces in $M_{i}$. Choose a subspace $V_{i}\subset H_{2}(M_{i})$ such that the map $H_{2}(M_{i})\to H_{2}(M_{i},\partial M_{i})$ restricts to an isomorphism between $V_{i}$ and $\overline{V_{i}}$. Identify $V_{i}$ with its image under the inclusion $H_{2}(M_{i})\subset H_{2}(M)$. Let $\mathcal{T}\subset H_{2}(M)$ be the subspace generated by the classes of JSJ tori of $M$. The null-space of the Thurston norm splits as the direct sum $\mathcal{T}\oplus V_{1}\oplus...\oplus V_{n}$. If we show that the dimension of $\mathcal{T}$ is $b_{1}(\Gamma)$, then the statement follows from Propositions 2.15 and 3.2.
@@
We first show that a class $\alpha\in H_{2}(M;\operatorname{\mathbb{Z}})$ is fibered if and only if $\psi_{i}(\alpha)\neq 0$ for each $i=1,...,n$. Suppose that $\psi_{i}(\alpha)\neq 0$ for each $i$. Without loss of generality, assume that $M$ is connected and $\alpha$ is primitive, so that $\alpha$ can be represented by a connected properly embedded oriented surface $S\subset M$. We can also assume that $S\cap M_{i}$ is a (possibly disconnected) horizontal surface in $M_{i}$. Cut $M$ open along $S$, then call $N$ the new manifold. Notice that the manifold $N_{i}$ obtained by cutting $M_{i}$ open along $S\cap M_{i}$ is an oriented interval-bundle over $S\cap M_{i}$, hence necessarily $(S\cap M_{i})\times[0,1]$. Properly speaking, $N_{i}$ is a manifold with corners, $\partial N_{i}$ being the union of the “vertical boundary” $(S\cap\partial M_{i})\times[0,1]$ and the “horizontal boundary” $(S\cap M_{i})\times\{0,1\}$. Since $\alpha\neq 0$, $S$ is non-separating and $N$ is the connected manifold obtained by gluing the interval bundles $(S\cap M_{i})\times[0,1]$ along their “vertical” boundary components. Each vertical boundary component is an interval-foliated annulus. Since the foliations of the annulus by intervals are all isotopic with each other, we can isotope the interval-bundles in a collar neighborhood of the vertical boundary of every $N_{i}$ so that they patch together to an interval-bundle of $N$ over the surface $S$. Again, this interval-bundle must be $S\times[0,1]$, thus $M$ fibers over the circle with fiber $S$.
For the vice-versa, suppose that $M$ fibers over the circle with fiber $S$. As in the proof of Proposition 3.2, $S$ cannot be a torus. In particular, since the fibration is a taut foliation, by [13] each JSJ torus of $M$ can be slightly isotoped to be transverse to the fibers of the fibration, apart for a finite number of saddle singularities. In fact, by the Poincare-Hopf Theorem, each JSJ torus can be taken to be entirely transverse to the fibration. Hence, the fibration restricts to a fibration onto every piece $M_{i}$. If any $S\cap M_{i}$ were vertical, then $M_{i}$ would be a torus-bundle or an annulus-bundle, which is excluded since $\chi_{i}<0$. It follows that $\psi_{i}([S])\neq 0$, for every $i=1,...,n$.
@@ 4 Characterization of the norms that can be realised
The formula (4) implies that the nonvanishing Thurston norm of a closed oriented graph manifold can be expressed as a sum of absolute values of linear functionals $H_{2}^{nv}(M)\to\operatorname{\mathbb{R}}$. Notice that those linear functionals are rational on $H_{2}^{nv}(M;\operatorname{\mathbb{Q}})$. In this section, we will prove that every norm with these properties can be realised as the nonvanishing Thurston norm on some closed oriented good graph manifold.
@@
Let |||||$\cdot$||| | ⋅ | | be a norm on $\operatorname{\mathbb{R}}^{d}$ that can be written as
@@
for some $\beta_{1},...,\beta_{n}\in\operatorname{\mathbb{Q}}^{d}$, and let $g_{1},...,g_{n}$ be nonnegative integers. There exist a closed oriented good graph manifold $M$ with base surfaces of genera $g_{1},...,g_{n}$ and a rational basis on $H_{2}(M)$, such that the nonvanishing Thurston norm on $H_{2}^{nv}(M)$ in the induced basis coincides with |||||$\cdot$||| | ⋅ | |. Moreover, we can ask for either all or none of the top-dimensional Thurston cones of $H_{2}(M)$ to be fibered.
@@
Finally, for the fibering statement, recall that a class $\alpha\in H_{2}(M)$ is fibered if and only if $\psi_{i}(\alpha)\neq 0$ for each $i=1,...,n$ (see the proof of Proposition 3.5). In particular, $M$ does not fiber over the circle if and only if there is an $i$ so that $\psi_{i}(\alpha)=0$ for every $\alpha\in H_{2}(M)$, i.e. $\beta_{i}=0$. Hence, if we want to construct an $M$ fibering over the circle, we just need to repeat the construction after discarding the $\beta_{i}$s which are zero. Oppositely, if we want $M$ not to fiber over the circle, then it is enough to repeat the construction with $\beta_{1},...,\beta_{n},\beta_{n+1}=0$. ∎
@@
The same reasoning as the one in [13, Chapter 4] shows that every such norm is a sum of absolute values of linear functionals $\operatorname{\mathbb{Q}}^{2}\to\operatorname{\mathbb{Q}}$. The statement follows from Theorem 4.1 by choosing all $g_{i}=0$. ∎
Unfortunately, we will see in the next section that the analogous statement does not hold in higher dimension.
@@ 5 Linear cellular decompositions of the sphere induced by Thurston norms
Now, we aim to explore geometrically the nonvanishing Thurston norms of graph manifolds. We will describe how to recover the shape of their unit polyhedra and observe a key property of them, that we will call “completeness”. Looking for completeness of the unit balls will allow us to show that the family of norms realisable as nonvanishing Thurston norms of graph manifolds is strictly contained in the family of nonvanishing Thurston norms of compact oriented $3$-manifolds (see Example 5.5).
@@
Let $P\subset\operatorname{\mathbb{R}}^{n}$ be a convex (solid) polyhedron containing the origin in its interior. Let $\pi:\operatorname{\mathbb{R}}^{n}-\{0\}\to S^{n-1}$ be the radial projection. The cellular decomposition of $\partial P$ into faces induces a cellular decomposition of $S^{n-1}$ via the projection $\pi$. If $F$ is a $k$-dimensional (open) face of $\partial P$, then $\pi(F)$ is a $k$-dimensional cell in $S^{n-1}$, contained in the intersection of $S^{n-1}$ with a $(k+1)$-dimensional subspace of $\operatorname{\mathbb{R}}^{n}$, for every $k\leq n-1$. We call such a cellular decomposition of $S^{n-1}$ a linear cellular decomposition and we indicate it with the symbol $\mathcal{F}(P)$. For $k=0,...,n-1$, $\mathcal{F}^{k}(P)$ will indicate the $k$-skeleton of $\mathcal{F}(P)$.
@@
If the inclusion above is an equality, we say that the linear cellular decomposition is complete, and the polyhedron $P$ is complete. When the $\beta_{i}$ are all in $\operatorname{\mathbb{Q}}^{n}$, we refer to the linear cellular decomposition as a rational cellular decomposition.
If |||||$\cdot$||| | ⋅ | | is a norm on $\operatorname{\mathbb{R}}^{n}$ whose unit ball is a polyhedron $P$, we call $\mathcal{F}(||\cdot||):=\mathcal{F}(P)$. Such a linear cellular decomposition is symmetric, meaning that $\mathcal{F}^{k}(||\cdot||)$ is preserved by the antipodal map $v\to-v$, for every $k=0,...,n-1$.
@@
Let $C$ be the cube $[-1,1]^{3}\subset\operatorname{\mathbb{R}}^{3}$, see Figure 2 left. The linear cellular decomposition of $S^{2}$ induced by $C$ is rational and symmetric but not complete, indeed
@@
The same line of reasoning as above shows that $v$ belongs to the open cone $C^{\prime}$ where $||\cdot||^{\prime}$ coincides with a linear function $L:\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}$. Since $C^{\prime}$ is open, there is an $\epsilon>0$ such that both $v+\epsilon\beta_{i}$ and $v-\epsilon\beta_{i}$ lie in $C^{\prime}$. We can then compute
@@
Let $L$ be a $3$-link chain and let $M:=S^{3}-\overset{\circ}{N}(L)$, see Figure 3. As shown in [13], the Thurston unit ball on $H_{2}(M,\partial M)$ is the polyhedron with vertices $(1,0,0),(0,1,0),(0,0,1),(-1,1,1)$ and their opposites. The linear cellular decomposition of $S^{2}$ induced by the Thurston norm is not complete, as there are vertices with odd degree. Hence, the family of norms that are realizable as nonvanishing Thurston norm of graph manifolds is strictly included in the family of norms realizable as nonvanishing Thurston norm of $3$-manifolds.
@@
In [13], Thurston showed that if a convex symmetric polygon in $\operatorname{\mathbb{R}}^{2}$ has vertices ±$\pm v_{1},...,\pm v_{k}$ in $\operatorname{\mathbb{Z}}^{2}$, with $v_{i}\equiv v_{j}$ modulo $2$ for all $i,j$, then it can be realised as the dual Thurston ball of a compact orientable $3$-manifold. As pointed out by Sane in [12], the idea at the base of Thurston’s proof is that every such polygon is the dual unit ball of an intersection norm on the first homology group of a torus. For every intersection norm on the first homology of a torus, Thurston finds a $3$-manifold whose Thurston norm on the second homology coincides with the given intersection norm. In particular, Sane generalizes this construction to the higher-dimensional case: he considers intersection norms on the first homology of the closed orientable surface of genus $g$ and their induced polyhedra in $\operatorname{\mathbb{R}}^{2g}$. As shown in [11], not every symmetric polyhedron is realizable this time, nor assuming the same parity condition working for polygons.
@@
It is worth stressing that both Corollaries 5.7 and 5.8 are here established with respect to the choice of a rational basis for the second homology of the relative graph manifold. In order to show the same results with respect to integral bases, it would be enough to show that for every $\beta_{1},...,\beta_{n}$ integral generators of $\operatorname{\mathbb{Z}}^{d}$ there exist a closed oriented good graph manifold $M$, an integral basis of $H_{2}(M)$ and a natural number $N$, such that the nonvanishing Thurston norm is $N\sum_{1}^{n}|\langle\beta_{i},\cdot\rangle|$ in the induced integral basis. Pursuing this strategy would demand a closer look to the relationship between $\psi(H_{2}^{nv}(M;\operatorname{\mathbb{Z}}))$ and $\ker A\cap\operatorname{\mathbb{Z}}^{n}$ (see Proposition 3.2).
@@
$\operatorname{Span}_{\operatorname{\mathbb{R}}}\{\beta_{j}\,|\;j\neq i\}$ has dimension $n-1$, then choose $\overline{\beta}_{i}$ to be a non null vector in $\bigcap_{j\neq i}\beta_{j}^{\bot}$, normalized to have $||\overline{\beta}_{i}||^{2}=|\langle\overline{\beta}_{i},\overline{\beta}_{i}\rangle|=1$. The unit ball of |||||$\cdot$||| | ⋅ | | is the polyhedron with vertices ±$\pm\overline{\beta}_{1},...,\pm\overline{\beta}_{n}$ in $\operatorname{\mathbb{R}}^{n}$.
@@
Identity (7) was proved by Gabai in [3]. In particular, the Thurston ball of $H_{2}(M)$ can be fully recovered by intersecting the Thurston ball of $H_{2}(\widetilde{M})$ with the subspace $\pi^{*}H_{2}(M)$ and rescaling it by a factor $2$.
@@
By Lemma 3.1 and Proposition 5.3, the nonvanishing Thurston unit ball on the nonvanishing second homology of a closed oriented graph manifold is a symmetric complete polyhedron, and it has rational vertices. Unfortunately, not every symmetric complete polyhedron with rational vertices can be obtained as the nonvanishing Thurston ball of some closed oriented graph manifold. As a counterexample, consider the complete polyhedron $P\subset\operatorname{\mathbb{R}}^{3}$ in Figure 5 with vertices